# DragText: Rethinking Text Embedding in Point-based Image Editing

### v1

@@ Abstract
Point-based image editing enables accurate and flexible control through content dragging. However, the role of text embedding in the editing process has not been thoroughly investigated. A significant aspect that remains unexplored is the interaction between text and image embeddings. In this study, we show that during the progressive editing of an input image in a diffusion model, the text embedding remains constant. As the image embedding increasingly diverges from its initial state, the discrepancy between the image and text embeddings presents a significant challenge. Moreover, we found that the text prompt significantly influences the dragging process, particularly in maintaining content integrity and achieving the desired manipulation. To utilize these insights, we propose DragText, which optimizes text embedding in conjunction with the dragging process to pair with the modified image embedding. Simultaneously, we regularize the text optimization process to preserve the integrity of the original text prompt. Our approach can be seamlessly integrated with existing diffusion-based drag methods with only a few lines of code.
Figure 1: In point-based image editing, a user first draws a mask on the image to define an editable region, and edits the image by “dragging” the contents from the user-defined handle points (red) to target points (blue). Our DragText stabilizes the point-based image editing process through parallel image and text optimization, consistently improving an array of existing diffusion-based drag methods.
*Equal contribution
Corresponding author
@@ 1 Introduction
Text-to-image (T2I) models have made significant advancements in image editing alongside the development of diffusion models [24, 26, 23, 28, 25, 1]. These models are effective in detailed modifications, such as inpainting[36, 34], style transfer [3, 39], and content replacement [27, 4, 9]. For instance, when a user inputs the text prompt "a dog looking up" into a T2I model, the resulting image shows the dog lifting its head. However, if one wishes to provide more explicit instructions for more detailed structural edits (e.g., movement angle or distance), designing appropriate text prompts to inject such a level of intention becomes far from trivial. To deal with this, several studies have shifted towards non-text controls (e.g., points, edges, poses, sketches) to avoid ambiguity and achieve controllability [21, 37, 35, 18]. Among these, point-based image editing (Fig. 1) is particularly noteworthy in that it employs pairs of instruction points, allowing fine-grained control.
@@
Point-based image editing has recently advanced with diffusion models, effectively manipulating both synthetic and real images across diverse domains [21, 29, 12, 13, 40]. To generalize image editing in various fields, large-scale pre-trained latent diffusion models have been utilized. These models are trained with cross-attention between text and image embeddings in denoising U-Net. Therefore, they also require text prompt as input. When point-based image editing drags content in the image embedding space, the text prompt provides conditions to transform the latent vector. However, to the best of our knowledge, there is no prior research which investigates the impact of text. Unlike T2I which actively analyzes text and is thus able to utilize text in innovative ways [3, 15, 4], the role of text and its potential in point-based image editing remains unknown.
@@
In this study, we explore how the text prompt affects dragging and discover whether the drag successfully moves to the desired position is influenced by the text prompt. As the image is optimized through drag editing, it naturally deviates from the original image in the image embedding space [19]. However, text embedding remains stationary and thus fails to accurately describe the edited image. This static text embedding is still used during the drag editing process and the denoising process in point-based image editing. For example, as shown in Fig. 2, in the editing process where a woman’s hair is made to move aside, the edited image is no longer coupled with the original text "A woman with a leaf tattoo on her neck." Therefore, as observed in the second image, it fails to reach the desired drag points, which we call drag halting, or it reaches the points but loses semantics.
@@
We propose DragText, a method designed to rectify the static text embedding in point-based image editing. DragText optimizes the original text embedding in a similar way as image optimization. “Dragging” text embedding in parallel with the image ensures that the edited text embedding remains coupled with the edited image embedding. As a result, drag-halting is alleviated, and the image embedding is sufficiently optimized and helps handle points reach their target points. Additionally, we designed a text embedding regularization technique to ensure the edited text embedding does not diverge too significantly from the original text embedding. It preserves key styles of the image while allowing for natural changes during drag editing.
@@
Figure 2: Illustration of the drag editing process within the image and text embedding spaces of the diffusion model (DM). During the edit, the original image embedding $\mathbf{z}_{t}$ naturally deviates to the dragged image latent vector $\mathbf{\bar{z}}_{t}$. With no text optimization, the corresponding text embedding $\mathbf{c}$ is decoupled from $\mathbf{\bar{z}}_{t}$. Hence, optimal text embedding $\mathbf{\hat{c}}$ coupled with dragged images has to be acquired to make the optimal latent vector $\mathbf{\hat{z}}_{t}$ which then holds the related semantics via text.
@@
Contributions. Our contributions are as follows: 1) We are the first to analyze the impact of text on dragging, a previously neglected yet crucial component of point-based image editing; 2) We propose DragText which optimizes text embedding that aids drag operations without forgetting content; 3) DragText can be easily applied in a plug-and-play manner to various diffusion-based drag models, consistently improving performance.
@@ 2.1 Point-based Image Editing
Task Definition. Point-based image editing enables precise modifications to images [2, 21, 16, 17]. As shown in Fig. 1, a user inputs the image to be edited, draws a mask representing an editable region, and specifies handle points along with their corresponding target locations. During the editing process, handle points are moved toward their respective target points through an image optimization approach. The positions of these points are continuously tracked and updated after each optimization step.
Existing Methods. DragGAN [21] is a representative method in point-based image editing that proposes latent code optimization and point tracking. However, it struggles with generalization, particularly with real images. DragDiffusion [29] employs a text-conditional diffusion model [26] to expand applicability and improve spatial control, following the method of DragGAN. Following these advancements, many point-based editing diffusion models have been developed. For instance, FreeDrag [12] uses original image features as a reference template and reduces the burden on point tracking with a line search. DragNoise [13] reduces computational load by optimizing the U-Net bottleneck feature instead of the latent vector. GoodDrag [40] alternates between dragging and denoising to prevent error accumulation and enhance image fidelity. As shown in these examples, previous works have primarily focused on image optimization to directly improve manipulation. Our approach emphasizes the underestimated role of the text prompt, offering a new perspective that enhances not only manipulation quality but also the dragging mechanism.
@@ 2.2 Text Optimization
Recent advancements in Vision-Language Models (VLMs) have significantly enhanced the ability to flexibly connect text and images [22, 8]. To improve these connections, various methods of text optimization are being explored. One approach is context optimization [42, 41], which refines the text used by pre-trained VLMs for a better understanding of handcrafted captions by models. Another method, developed by Gal et al. [3], optimizes a single word to effectively convey content information, thereby personalizing text-based image editing. Additionally, Mokady et al. [15] introduce a technique for optimizing null-text "", addressing the issue of images deviating from their original trajectory due to accumulated errors from classifier-free guidance [6]. Our work is inspired by these prior studies which closely examine the strong coupling between text and image.
@@ 3 Motivations
Editing an image causes the original image embedding to move in the image embedding space. Therefore, the edited image embedding must follow a different denoising trajectory compared to the original image embedding [19, 15]. As text and image interact through cross-attention, we hypothesize that text embedding also plays a crucial role in determining this trajectory.
@@ 3.1 Text Prompt in Diffusion Model
Role in Inversion and Denoising Processes. The inversion process pairs the image embedding with text embedding, regardless of whether the text embedding semantically matches the image. Maintaining this pairing during the denoising process is important for the fidelity of image sampling.
In Fig. 3 (a), we examine differences in image sampling outcomes based on whether the text embedding is also paired in the denoising process. Our analysis reveals that using the paired text embedding during the denoising process enables the accurate sampling of the original image. However, if unpaired text embedding is employed, the image’s style alters, and the model fails to sample the original image accurately. This observation highlights the necessity of text embedding-image embedding pairing to maintain the integrity of the image’s attributes. Moreover, it raises concerns about the appropriateness of using the original input text embedding in the denoising process of dragging, as the image is modified during the drag editing process, it can no longer be accurately paired with the text embedding.
Remark. To achieve the appropriate image sampling result, the text embedding used after inversion remains consistently paired with the image embedding.
Figure 3: Red boxes in (a) imply the consistent use of text prompt across diffusion processes, resulting in high-fidelity sampling. During dragging the image in (b), the original text prompt and its alternative intention text prompt, crafted from prompt engineering, are insufficient to prevent drag halting.
@@ 3.2 Text Embedding in Point-based Image Editing
Challenges in Original Text. During dragging, the image embedding gradually changes to reflect the manipulation. However, the text embedding remains static, restricting the extent of image manipulation.
As shown in the top-left image of Fig. 3 (b), the original text embedding fails to reach the target point during the dragging process. Moreover, since the original text embedding is paired with the input image, the edited image embedding does not maintain a strong coupling with the text embedding. To avoid this weakly-coupling issue, one might hypothesize that this issue would be resolved with unconditional text embedding (i.e. null text). However, the top-right image of Fig. 3 (b) shows that the model fails to maintain the semantic information of the object without text conditions. Therefore, we search for an alternative text embedding that mitigates drag halting while preserving semantics.
Remark. While text embedding plays a crucial role in maintaining semantics, it can also impede the handle points from reaching the target points.
Figure 4: The pipeline of DragText. The image $\mathbf{x}_{0}$ is mapped to a low-dimensional space through a VAE encoder, and the text is encoded by a CLIP text encoder as the text embedding $\mathbf{c}$. Through DDIM inversion with $\mathbf{c}$, the latent vector $\mathbf{z}_{t}$ is obtained. At time step $t=35$, $\mathbf{z}^{0}_{t}$ and $\mathbf{c}$ are optimized to $\mathbf{\hat{z}}^{k}_{t}$ and $\mathbf{\hat{c}}$ by iterating with motion supervision (M.S.) and point tracking (P.T.) $k$-times.
@@
To validate its effectiveness, we craft intention text prompt that describes the user’s intention to edit the image, such as "a photo of a short neck jug and a glass", and apply it during the denoising process, as illustrated in Fig. 3. However, this approach [14] encounters the same issue as the original prompt due to the weak coupling between image and text embeddings during the dragging process (b). Thus, we linearly interpolate the intention text embedding with the original text embedding during drag editing. Nevertheless, actively reflecting the gradual changes in the image embedding remains limited. Consequently, we aimed to develop a more sophisticated text embedding modification method in parallel with the gradual image manipulations.
@@
Necessity for Text Embedding Optimization. Based on this series of findings, we advocate for the necessity of alternative text embedding that can: 1) successfully reach the target point while maintaining semantics; 2) by gradually reflect changes introduced during image editing; 3) thus maintain pairing with the edited image embedding to ensure proper denoising. Furthermore, rather than relying on heuristic and highly uncertain methods such as prompt engineering, we propose an optimization approach that is intrinsically parallel to image editing and can be strongly coupled with the edited image.
@@ 4 Method
In this study, we propose DragText, a novel framework to optimize text embedding along with point-based image editing. We briefly explain the diffusion process first (Sec. 4.1), and then describe the drag editing process in detail, including text embedding optimization (Sec. 4.2). Representative diffusion model-based dragging methods [29, 13, 12, 40] share the same approach, therfore, we explain DragText based on DragDiffusion [29], and minor modifications for others are in the Supplement.
@@ 4.1 Preliminaries on Diffusion Models
Diffusion models generate images through a forward process and a reverse process. By using a denoising U-Net, these models predict noise at a specific time-step $t$ and progressively denoise it to create an image. Unlike DDPM [5], DDIM [32] allows for deterministic image generation, which has made real image editing possible.
Recent advancements in Latent Diffusion Models (LDM) [26] have enabled the diffusion processes to occur within the latent space. The input image $\mathbf{x}_{0}$ is encoded with VAE encoder [10] to make latent vector $\mathbf{z}_{0}$. LDM converts $\mathbf{z}_{0}$ into a noisy latent vector $\mathbf{z}_{t}$ using a forward process, allowing for image editing by manipulating this latent vector. Additionally, cross-attention mechanisms allow for conditioning with different prompts, making text-based conditioning feasible with text embedding $\mathbf{c}$. The $\mathbf{z}_{t}$ is denoised using the predicted $\mathbf{\epsilon_{\theta}}(\mathbf{z}_{t},t,\mathbf{c})$ from the denoising U-Net and then converted into an image $\mathbf{\hat{x}_{0}}$ through a VAE decoder. In our work, we use a pre-trained LDM Stable Diffusion [26], which is employed in previous studies.
@@ 4.2 Drag Editing
As shown in Fig. 4, we optimize the latent vector $\mathbf{z}_{t}$ and the text embedding $\mathbf{c}$ according to $n$ number of instruction points $\{h_{i}=(x_{i},y_{i}),\ g_{i}=(\tilde{x}_{i},\ \tilde{y}_{i}):i=1,2,\ldots,n\}$ via three stages: motion supervision, text embedding optimization, and point tracking. These stages are sequentially repeated until all handle points have reached corresponding target points or the maximum number of iteration steps $k$ is achieved.
Figure 5: Comparison of editing results with and without applying DragText to the baseline model. DragText shows superior semantic control and dragging.

### v2

@@ Abstract
Point-based image editing enables accurate and flexible control through content dragging. However, the role of text embedding during the editing process has not been thoroughly investigated. A significant aspect that remains unexplored is the interaction between text and image embeddings. During the progressive editing in a diffusion model, the text embedding remains constant. As the image embedding increasingly diverges from its initial state, the discrepancy between the image and text embeddings presents a significant challenge. In this study, we found that the text prompt significantly influences the dragging process, particularly in maintaining content integrity and achieving the desired manipulation. Upon these insights, we propose DragText, which optimizes text embedding in conjunction with the dragging process to pair with the modified image embedding. Simultaneously, we regularize the text optimization process to preserve the integrity of the original text prompt. Our approach can be seamlessly integrated with existing diffusion-based drag methods, enhancing performance with only a few lines of code.
Figure 1: In point-based image editing, a user first draws a mask on the image to define an editable region, then edits the image by “dragging” the contents from the user-defined handle points (red) to target points (blue). Our DragText stabilizes the process of point-based image editing through parallel image and text optimization, consistently improving an array of existing diffusion-based drag methods.
*Corresponding author
@@ 1 Introduction
Text-to-image (T2I) models have made significant advancements in image editing alongside the development of diffusion models [26, 28, 25, 30, 27, 1]. These models are effective in broad modifications, such as inpainting[38, 36], style transfer [3, 41], and content replacement [29, 5, 10]. For instance, when a user inputs the text prompt “a dog looking up” into a T2I model, the resulting image shows the dog lifting its head. However, if one wishes to provide clearer explicit instructions for more detailed structural edits (e.g., movement angle or distance), designing appropriate text prompts to inject such a level of intention becomes far from trivial. To deal with this, several studies have shifted towards non-text controls (e.g., points, edges, poses, sketches) to avoid ambiguity and achieve controllability [22, 39, 37, 19]. Among these, point-based image editing (Fig. 1) is particularly noteworthy in that it employs pairs of instruction points, allowing fine-grained control.
@@
Point-based image editing has recently advanced with diffusion models, effectively manipulating both synthetic and real images across diverse domains [22, 31, 13, 14, 42]. To generalize image editing in various fields, large-scale pre-trained latent diffusion models have been utilized. These models are trained with cross-attention between text and image embeddings in denoising U-Net. Therefore, they also require text prompts as input. When point-based image editing drags content in the image embedding space, the text prompt provides conditions to the latent vector. However, to the best of our knowledge, no prior research has investigated the impact of text. Unlike T2I methods which actively analyze text and utilize text in innovative ways [3, 16, 5, 4, 23], the role of text and its potential in point-based image editing remains unknown.
@@
In this study, we explore how the text prompt affects the dragging process and discover whether the content successfully dragged to the desired position is influenced by the text prompt. As the image is optimized through drag editing, it naturally deviates from the original image in the image embedding space [20]. However, the text embedding remains stationary and thus fails to describe the edited image accurately. This static text embedding is used during the drag editing process and the denoising process in point-based image editing [31, 13, 14, 42], leading to what we term drag halting, where the drag editing process fails to reach the intended drag points, or it reaches them but with a loss of semantic integrity. For example, as shown in Fig. 2, during the editing process of moving a woman’s hair aside, the edited image is no longer strongly coupled with the original text “A woman with a leaf tattoo on her neck.”. Despite the change in the woman’s hair, the original text prompt does not reflect this change. As a result, the handle point falls short of fully reaching the target points as seen in the middle image.
@@
We propose DragText, a method designed to rectify the static text embedding in point-based image editing. DragText optimizes the original text embedding in a similar way to image optimization. “Dragging” text embedding in parallel with the image ensures that the edited text embedding remains coupled with the edited image embedding. It alleviates drag halting by optimizing the image embedding, which helps handle points reach their targets. Moreover, we designed a text embedding regularization technique to ensure the edited text embedding does not diverge too significantly from the original text embedding. It preserves key styles of the image while allowing for natural changes within the image during drag editing.
@@
Figure 2: Illustration of the drag editing process within the image and text embedding spaces of the diffusion model (DM). During editing, the original image embedding $\mathbf{z}_{t}$ naturally deviates to the dragged image latent vector $\mathbf{\bar{z}}_{t}$. Without text optimization, the corresponding text embedding $\mathbf{c}$ is decoupled from $\mathbf{\bar{z}}_{t}$, resulting in drag halting. Hence, optimal text embedding $\mathbf{\hat{c}}$ coupled with dragged images has to be acquired to make the optimal latent vector $\mathbf{\hat{z}}_{t}$ which then holds the related semantics via text.
@@
Contributions. Our contributions are as follows: 1) We are the first to analyze the impact of text prompts on the dragging process, a previously neglected yet crucial component of point-based image editing; 2) We propose DragText which optimizes text embedding that aids drag operations without forgetting content; 3) DragText can be easily applied in a plug-and-play manner to various diffusion-based point-based image editing methods, consistently improving performance.
@@ 2.1 Point-based Image Editing
Task Definition. Point-based image editing enables precise modifications to images via point instruction [2, 22, 17, 18]. As shown in Fig. 1, a user inputs the image to edit, draws a mask representing an editable region, and specifies handle points along with their corresponding target locations. During the editing process, handle points are moved toward their respective target points through an image optimization strategy. The positions of these points are continuously tracked and updated after each optimization step.
Existing Methods. DragGAN [22] is a representative method in point-based image editing that proposes latent code optimization and point tracking. However, it struggles with generalization, particularly with real images. DragDiffusion [31] employs a text-conditional diffusion model [28] to expand applicability and improve spatial control, following the strategy of DragGAN. Several point-based editing diffusion methods have been developed upon these advancements. For instance, FreeDrag [13] uses original image features as a reference template and reduces the complexity of point tracking with a line search. DragNoise [14] reduces computational load by optimizing the U-Net bottleneck feature instead of the latent vector. GoodDrag [42] alternates between dragging and denoising processes to prevent error accumulation and enhance image fidelity. Previous works have primarily focused on image optimization to directly improve manipulation. Unlike these, our approach emphasizes the underestimated role of the text prompt, providing a new perspective of enhancing not only manipulation quality but also the dragging mechanism.
@@ 2.2 Text Optimization
Recent advancements in Vision-Language Models (VLMs) have significantly enhanced the ability to flexibly connect text and images [24, 9]. Various methods of text optimization are being explored to improve these connections. One approach is context optimization [44, 43], which refines text prompts to help VLMs better understand hand-crafted captions. Another method, developed by Gal et al. [3], optimizes a single word to effectively convey content information, thereby personalizing text-based image editing. Mokady et al. [16] introduces the strategy to optimize null-text “”, which addresses the issue of images deviating from their original trajectory due to accumulated errors from classifier-free guidance [7]. Our work is inspired by these prior studies which closely examine the strong coupling between text and image. However, our approach differs from these in that we perform text optimization simultaneously with image optimization in point-based image editing to avoid increasing the optimization steps and time.
@@ 3 Motivations
Editing an image causes the original image embedding to move in the image embedding space. Therefore, the edited image embedding must follow the denoising trajectory that differs from that of the original image embedding [20, 16]. We hypothesize that text embedding also plays a crucial role in determining this distinct trajectory as text and image interact through cross-attention.
@@ 3.1 Text Prompt in Diffusion Model
Role in Inversion and Denoising Processes. The inversion process pairs the image embedding with text embedding, regardless of whether the text embedding semantically matches the image. Maintaining this pairing in the denoising process is important for the fidelity of image sampling.
In Fig. 3 (a), we examine differences in image sampling outcomes based on whether the text embedding is also paired in the denoising process. Our analysis reveals that using the paired text embedding in the denoising process (red boxes) enables the accurate sampling of the original image. However, if unpaired text embedding is employed, the image’s style alters, and the model fails to sample the image accurately. This observation highlights the necessity of text-image embedding pairing to maintain the integrity of the image’s attributes. Moreover, it raises concerns about the appropriateness of using the original text embedding in the denoising process of point-based image editing, as the image is modified during dragging, it can no longer be accurately paired with the original text embedding.
Remark. To achieve the appropriate image sampling result, the text embedding used after the inversion process remains consistently paired with the image embedding.
Figure 3: (a) Red boxes imply the consistent use of text prompts across diffusion processes, resulting in high-fidelity sampling. In contrast, other images result from inconsistent text prompt usage, leading to inaccurate sampling outcomes. (b) The original text prompt and its alternative intention text prompt, crafted from prompt engineering, are insufficient to prevent drag halting.
Figure 4: The pipeline of DragText. The image $\mathbf{x}_{0}$ is encoded through a VAE encoder as the latent vector $\mathbf{z}$, and the text is encoded by a CLIP text encoder as the text embedding $\mathbf{c}$. Through DDIM inversion with $\mathbf{c}$, the latent vector $\mathbf{z}_{t}$ is obtained. At time step $t=35$, $\mathbf{z}^{0}_{t}$ and $\mathbf{c}$ are optimized to $\mathbf{\hat{z}}^{k}_{t}$ and $\mathbf{\hat{c}}$ by iterating with motion supervision (M.S.), text optimization, and point tracking (P.T.) $k$-times.
@@ 3.2 Text Embedding in Point-based Image Editing
Challenges in Original Text. During dragging, the image embedding incrementally adapts to reflect the manipulation. However, the text embedding remains static, limiting the extent of image manipulation.
As shown in the top-left image of Fig. 3 (b), the original text embedding fails to reach the target point during the dragging process. Since the original text embedding is paired with the input image, the edited image embedding does not maintain a strong coupling with the text embedding. To avoid this weakly-coupling issue, one might hypothesize that unconditional text embedding (i.e. null text) could resolve the problem. However, the top-right image of Fig. 3 (b) shows that the model fails to maintain the semantic information of the object without text conditions. Therefore, we search for an alternative text embedding that mitigates drag halting while preserving semantics.
Remark. While text embedding plays a crucial role in maintaining semantics it can also impede the handle points from reaching the target points.
@@
To validate the effectiveness of prompt engineering [15], we craft an intention text prompt that describes the user’s intention to edit the image, such as “a photo of a short neck jug and a glass”, and apply it during the denoising process, as illustrated in Fig. 3. However, this approach encounters the same issue as the original prompt due to the weak coupling between image and text embeddings. Thus, we linearly interpolate the intention text embedding with the original text embedding during drag editing. Nevertheless, actively reflecting the gradual changes in the image embedding remains limited. Consequently, we aimed to develop a more sophisticated text embedding modification method in parallel with the gradual image manipulations. A detailed description of the methods and results of this analysis can be found in Appendix B.
@@
Necessity for Text Embedding Optimization. Based on this series of findings, we advocate for the necessity of alternative text embedding that can: 1) successfully reach the target point while maintaining semantics; 2) by gradually reflect changes introduced during image editing; 3) thus maintain pairing with the edited image embedding to ensure proper denoising. Furthermore, rather than relying on heuristic and inconsistent methods such as prompt engineering, we propose an optimization approach that is intrinsically parallel to image editing and can be strongly coupled with the edited image.
@@ 4 Method
In this study, we propose DragText, a novel framework to optimize text embedding along with point-based image editing. We briefly explain the diffusion process first (Sec. 4.1), and then describe the drag process in detail, including text embedding optimization (Sec. 4.2). Representative diffusion model-based dragging methods [31, 14, 13, 42] share the same approach for motion supervision and point tracking based on DragGAN [22] and DragDiffusion [31]. Therefore, we explain DragText primarily with reference to DragDiffusion [31], with minor modifications for others detailed in Appendix C.
@@ 4.1 Preliminaries on Diffusion Models
Diffusion models generate images through a forward process and a reverse process. By using a denoising U-Net, these models predict noise at a specific time-step $t$ and progressively denoise it to create an image. Unlike DDPM [6], DDIM [34] allows for deterministic image generation, making real image editing possible.
Recent developments in Latent Diffusion Models (LDM) [28] have advanced the diffusion processes to occur efficiently within the latent space. The input image $\mathbf{x}_{0}$ is encoded via VAE encoder [11] to form a latent vector $\mathbf{z}_{0}$, which then undergoes a forward process to transition into a noisy latent vector $\mathbf{z}_{t}$. This transformation facilitates image editing through the manipulation of this latent vector. Furthermore, cross-attention mechanisms provide conditioning capabilities with diverse prompts, making text-based conditioning feasible with text embedding $\mathbf{c}$. The $\mathbf{z}_{t}$ is subsequently denoised by using the predicted noise $\mathbf{\epsilon_{\theta}}(\mathbf{z}_{t},t,\mathbf{c})$ from a denoising U-Net, and finally, VAE decoder reconstructs the denoised latent vector into an image $\hat{\mathbf{x}}_{0}$. In our work, we use the pre-trained LDM Stable Diffusion [28], which is employed in previous studies [31, 14, 13, 42].
@@ 4.2 Drag Editing
As shown in Fig. 4, we optimize the latent vector $\mathbf{z}_{t}$ and the text embedding $\mathbf{c}$ according to $n$ number of instruction points $\{h_{i}=(x_{i},y_{i}),\ g_{i}=(\tilde{x}_{i},\ \tilde{y}_{i}):i=1,2,\ldots,n\}$ via three stages: motion supervision, text optimization, and point tracking. These stages are sequentially repeated until all handle points have reached corresponding target points or the maximum number of iteration steps $K$ is achieved.
Figure 5: Qualitative results of DragText. As demonstrated in the comparison of editing results with and without the application of DragText to the baseline model, DragDiffusion [31] integrating DragText demonstrates improved semantic control and precision in dragging.